# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ 2.1 Notation
In the following we think of states as non-negative linear functionals from a $C^{*}$-algebra to $\mathbb{C}$. For finite dimensional systems, we can simply think of a density matrix $\rho\in\mathbb{M}_{n}$ as a functional on $\mathbb{M}_{n}$, $\rho(X)=\operatorname{Tr}[\rho X]$. Let us fix the notations for states on an infinite chain following [8]. We consider states on an infinite chain with sites labelled by $n\in\mathbb{Z}$, where to each site $n$ a finite dimensional $C^{*}$-algebra $\mathcal{A}_{n}\equiv\mathcal{A}$ is associated. Without loss of generality, $\mathcal{A}$ can be embedded into the algebra of $d_{\mathcal{A}}\times d_{\mathcal{A}}$ matrices $\mathbb{M}_{d_{\mathcal{A}}}$. For each finite subset $X$ of $\mathbb{Z}$ we can define an algebra $\mathcal{A}_{X}:=\bigotimes_{n\in X}\mathcal{A}_{n}$. The algebra of infinite subsets is defined through the inductive limit, with the identifications $\mathcal{A}_{X_{1}}\equiv\mathcal{A}_{X_{1}}\otimes\mathds{1}_{X_{2}\setminus X_{1}}$ for $X_{1}\subseteq X_{2}$ finite subsets. The chain algebra $\mathcal{A}_{\mathbb{Z}}$ is defined in this way, for example, as the inductive limit for sets $[-n,n]:=\{-n,\dots,n\}$. The right chain algebra $\mathcal{A}_{R}$ is defined as the inductive limit of $\mathcal{A}_{[1,n]}$, and the left chain algebra $\mathcal{A}_{L}$ is defined as the inductive limit of $\mathcal{A}_{[-n,0]}$. The translation operators $\alpha_{r}$ act on the algebra by sending $\mathcal{A}_{X}$ into $\mathcal{A}_{X+r}$. We denote the set of translation invariant states by $\mathcal{T}$.
@@
Additionally, we use the following notation for linear maps. We denote by $\|T\|_{p}=(\operatorname{Tr}[(T^{\dagger}T)^{{p}/{2}}])^{1/p}$ the Schatten $p$-norm for $T:H_{1}\rightarrow H_{2}$ being a linear map between finite dimensional Hilbert spaces $H_{1},\,H_{2}$, where $T^{\dagger}$ is the adjoint map (we also use the notation $U^{\intercal}$ if we deal with maps between real vector spaces). The singular values of $T$ are denoted and ordered as $\sigma_{1}(T)\geq\sigma_{2}(T)\geq\dots\geq\sigma_{r}(T)$, so that $\|T\|_{p}=\left(\sum_{i=1}^{r}\sigma_{i}(T)^{p}\right)^{1/p}$. We denote by $T^{+}$ the Moore-Penrose pseudoinverse of $T$. For a map $L:B_{1}\rightarrow B_{2}$, where $B_{1}$ and $B_{2}$ are spaces of finite dimensional linear maps equipped with Schatten norms, we have $\|L\|_{p\rightarrow p}:=\sup_{X\in B_{1},\|X\|_{p}\leq 1}\|L(X)\|_{p}$, and $\|L\|_{p\rightarrow p,cb}:=\sup_{d\in\mathbb{N}}\sup_{X\in\mathbb{M}_{d}\otimes B_{1},\|X\|_{p}\leq 1}\|(\mathrm{id}_{\mathbb{M}_{d}}\otimes L)(X)\|_{p}$. By slight abuse of notations, for $p\in[1,\infty]$, we will also denote by $\|\cdot\|_{p}$ the usual $\ell_{p}$ norm on $\mathbb{R}^{m}$, as well as the $p$-norm of linear functionals: given a functional $\varphi$ on a (matrix) space $\mathcal{C}$, $\|\varphi\|_{p}:=\sup_{\|X\|_{p^{\prime}}\leq 1}|\varphi(X)|$, where $11/p^{\prime}+1/p=$ 1. For $p=2$, this is also equal to $\|\varphi\|_{2}=(\sum_{i}|\varphi(X_{i})|^{2})^{1/2}$ for some orthonormal basis $\{X_{i}\}$ of $\mathcal{C}$.
2.2 Finitely correlated states
@@
Furthermore, also as part of [8, Prop. 2.1] we have
Proposition 2.3.
Realizations of minimal dimension, called regular realizations, are determined up to linear isomorphisms: if $(V,e,\mathbb{E},\rho)$ and $(V^{\prime},e^{\prime},\mathbb{E}^{\prime},\rho^{\prime})$ are two regular realizations of $\omega$, then $V$ and $V^{\prime}$ are isomorphic via a unique invertible linear map $M:V\rightarrow V^{\prime}$, such that $e^{\prime}=Me$, $\rho^{\prime}=\rho M^{-1}$, $\mathbb{E}^{\prime}_{A}=M\mathbb{E}_{A}M^{-1}$.
In the proof of the above statements, it is used that a realization of minimal dimension can be obtained as a quotient realization of any realization. Note, that if $\tilde{W}\subset V^{*}$ is a set of linear functionals, we denote the subspace on which they all vanish simultaneously by $\tilde{W}^{\perp}$ (that is, $\tilde{W}^{\perp}$ is just the intersection of their kernels).
Proposition 2.4.
@@
2.2.1 Observable regular realization
The proof of Proposition 2.2 gives also a way to construct a realization from the knowledge of a marginal of the state on a finite subset of the chain. We call this realization observable regular realization, as it is expressed only in terms of quantities which can be estimated from experiments.
@@ 2.2.1 Observable regular realization
with $X\in\mathcal{C}_{R}$, has dimension $m$. We call this vector space $V$, which can be thought of as a subspace of ${\mathcal{A}_{\Lambda}^{*}\cong\mathcal{A}_{\Lambda}}$, for some finite subset $\Lambda$ of the chain, say $\Lambda=[1,t^{*}]$ for some $t^{*}\leq m$ (see Appendix D), and define $L$ to be a projection from $\mathcal{A}_{\Lambda}^{*}$ to $V$.
@@
Proposition 2.5 (Observable regular realization).
@@
Let $(V,u,\mathbb{E},\sigma)$ be any realization and $(V^{\prime},e^{\prime},\mathbb{K}^{\prime},\rho^{\prime})$ its quotient realization. Since $(V^{\prime},e^{\prime},\mathbb{K}^{\prime},\rho^{\prime})$ and the observable realization $(\mathbb{C}^{m},e,\mathbb{K},\rho)$ constructed in Proposition 2.5 are both regular, there is a unique invertible linear transformation $M:V^{\prime}\rightarrow\mathbb{C}^{m}$ such that
@@
A specific map $M$ with desirable properties can be obtained explicitly as follows: Recall that $\mathcal{C}_{R}\subseteq\mathcal{A}_{[1,t^{*}]}$ and $\mathcal{C}_{L}\subseteq\mathcal{A}_{[1-t^{*},0]}$ (there is always such $t^{*}$, in fact $t^{*}\leq m$, see Appendix D). We define the map $\mathcal{F}:V\rightarrow\mathcal{A}_{[1-t^{*},0]}^{*}$ as
@@
Note, that $\mathcal{F}\mathbb{E}_{A}=\Psi_{A}$ and thus $\ker(U^{\intercal})=\mathcal{F}(W)^{\perp}$. Now, $Z\in\ker(\mathcal{F})$ if and only if for each $A$ we have
@@
In particular, $(\mathbb{C}^{m},e,\mathbb{K},\rho)$ is similar to the quotient realization of $(V^{\prime},e^{\prime},\mathbb{E},\rho^{\prime})$ through the invertible map $M:V^{\prime}\rightarrow\mathbb{C}^{m}$ defined by $ML=U^{\intercal}\mathcal{F}|_{W}$ (see Proposition 2.4).
@@
It holds that $\mathcal{F}(u)=\Omega(\mathds{1})$ so $U^{\intercal}\mathcal{F}(u)=e$ is immediate. Next, note that by definition $\rho U^{\intercal}\mathcal{F}=\tau\mathcal{F}$. Hence, if $w=\mathbb{E}_{A}(u)\in W$ then
@@
Writing $U^{\intercal}\mathcal{F}=ML$ by Lemma 2.6 the proof is complete. ∎
2.2.2 $C^{*}$-realizations
@@
We use the symbol $\mathcal{E}$ instead of $\mathbb{E}$ when dealing with realizations to distinguish them among realizations. Note, that whenever we find a realization with memory system being a $C^{*}$-algebra and generating map $\mathbb{E}$ being completely positive in the natural $C^{*}$ order, we can always find a realization satisfying the same that is of the form described in Definition 2.8, even with $\rho_{0}$ being of full support, see [8, Lem. 2.5].
As before we will also use the notation $\mathcal{E}_{A}(B)=\mathcal{E}(A\otimes B)$.
We can apply the construction of the quotient realization Proposition 2.4, that is we set
@@
Lemma 2.6 then takes the following form:
Lemma 2.9.
@@
2.2.3 Empirical realizations
It will be useful to also consider the following realizations, obtained by a similarity transformation of the observable realization. In the following $\hat{U}$ can be thought as a truncated left orthogonal map in the SVD of $\hat{\Omega}$, which is an estimate of $\Omega$ (cf.  Section 4). While the realization constructed from $\hat{\Omega}$ will not be an exact realization of $\omega$, the empirical realization constructed here is, and will be easier to compare with the estimate than the observable realization.
Proposition 2.10 (Empirical realizations).
@@ 2.2.3 Empirical realizations
where $M$ is the linear map connecting a quotient realization with the observable realization, as in Proposition 2.7. The following proposition easily follows.
Proposition 2.11 (Empirical realizations and quotient realization).
For an empirical realization as in Proposition 2.10, and a quotient realization as in Proposition 2.7, we have
@@
$\tilde{\mathbb{K}}_{A}\hat{M}L$ $=M\mathbb{K}^{\prime}_{A}L=ML\mathbb{E}_{A}|_{W},$ (47)
@@
3 Operator systems and realizations

### v3

@@ A.1 Notation
In the following we think of states as non-negative linear functionals from a $C^{*}$-algebra to $\mathbb{C}$. For finite dimensional systems, we can simply think of a density matrix $\rho\in\mathbb{M}_{n}$ as a functional on $\mathbb{M}_{n}$, $\rho(X)=\operatorname{Tr}[\rho X]$. Let us fix the notations for states on an infinite chain following [14]. We consider states on an infinite chain with sites labelled by $n\in\mathbb{Z}$, where to each site $n$ a finite dimensional $C^{*}$-algebra $\mathcal{A}_{n}\equiv\mathcal{A}$ is associated. Without loss of generality, $\mathcal{A}$ can be embedded into the algebra of $d_{\mathcal{A}}\times d_{\mathcal{A}}$ matrices $\mathbb{M}_{d_{\mathcal{A}}}$. For each finite subset $X$ of $\mathbb{Z}$ we can define an algebra $\mathcal{A}_{X}:=\bigotimes_{n\in X}\mathcal{A}_{n}$. The algebra of infinite subsets is defined through the inductive limit, with the identifications $\mathcal{A}_{X_{2}}\equiv\mathcal{A}_{X_{1}}\otimes\mathds{1}_{X_{2}\setminus X_{1}}$ for $X_{1}\subseteq X_{2}$ finite subsets. The chain algebra $\mathcal{A}_{\mathbb{Z}}$ is defined in this way, for example, as the inductive limit for sets $[-n,n]:=\{-n,\dots,n\}$. The right chain algebra $\mathcal{A}_{R}$ is defined as the inductive limit of $\mathcal{A}_{[1,n]}$, and the left chain algebra $\mathcal{A}_{L}$ is defined as the inductive limit of $\mathcal{A}_{[-n,0]}$. The translation operators $\alpha_{r}$ act on the algebra by sending $\mathcal{A}_{X}$ into $\mathcal{A}_{X+r}$. We denote the set of translation invariant states by $\mathcal{T}$.
@@
Additionally, we use the following notation for linear maps. We denote by $\|T\|_{p}=(\operatorname{Tr}[(T^{\dagger}T)^{{p}/{2}}])^{1/p}$ the Schatten $p$-norm for $T:H_{1}\rightarrow H_{2}$ being a linear map between finite dimensional Hilbert spaces $H_{1},\,H_{2}$, where $T^{\dagger}$ is the adjoint map (we also use the notation $T^{\intercal}$ if we deal with real maps). The singular values of $T$ are denoted and ordered as $\sigma_{1}(T)\geq\sigma_{2}(T)\geq\dots\geq\sigma_{r}(T)$, so that $\|T\|_{p}=\left(\sum_{i=1}^{r}\sigma_{i}(T)^{p}\right)^{1/p}$. We denote by $T^{+}$ the Moore-Penrose pseudoinverse of $T$. For a map $L:B_{1}\rightarrow B_{2}$, where $B_{1}$ and $B_{2}$ are spaces of finite dimensional linear maps equipped with Schatten norms, we have $\|L\|_{p\rightarrow p}:=\sup_{X\in B_{1},\|X\|_{p}\leq 1}\|L(X)\|_{p}$, and $\|L\|_{p\rightarrow p,cb}:=\sup_{d\in\mathbb{N}}\sup_{X\in\mathbb{M}_{d}\otimes B_{1},\|X\|_{p}\leq 1}\|(\mathrm{id}_{\mathbb{M}_{d}}\otimes L)(X)\|_{p}$. By slight abuse of notations, for $p\in[1,\infty]$, we will also denote by $\|\cdot\|_{p}$ the usual $\ell_{p}$ norm on $\mathbb{R}^{m}$, as well as the $p$-norm of linear functionals: given a functional $\varphi$ on a (matrix) space $\mathcal{C}$, $\|\varphi\|_{p}:=\sup_{\|X\|_{p^{\prime}}\leq 1}|\varphi(X)|$, where $11/p^{\prime}+1/p=$ 1. For $p=2$, this is also equal to $\|\varphi\|_{2}=(\sum_{i}|\varphi(X_{i})|^{2})^{1/2}$ for some orthonormal basis $\{X_{i}\}$ of $\mathcal{C}$.
A.2 Finitely correlated states
@@
Furthermore, also as part of [14, Prop. 2.1] we have
Proposition A.3.
Realizations of minimal dimension, called regular realizations, are determined up to linear isomorphisms: if $(V,e,\mathbb{E},\rho)$ and $(V^{\prime},e^{\prime},\mathbb{E}^{\prime},\rho^{\prime})$ are two regular realizations of $\omega$, then $V$ and $V^{\prime}$ are isomorphic via an invertible linear map $M:V\rightarrow V^{\prime}$, such that $e^{\prime}=Me$, $\rho^{\prime}=\rho M^{-1}$, $\mathbb{E}^{\prime}_{A}=M\mathbb{E}_{A}M^{-1}$.
In the proof of the above statements, it is used that a realization of minimal dimension can be obtained as a quotient realization of any realization. Note, that if $\tilde{W}\subset V^{*}$ is a set of linear functionals on $V$, we denote the subspace on which they all vanish simultaneously by $\tilde{W}^{\perp}$ (that is, $\tilde{W}^{\perp}$ is just the intersection of their kernels). Then we hsve the following.
Proposition A.4.
@@
A.2.1 Observable regular realization
The proof of Proposition A.2 gives also a way to construct a realization from the knowledge of a marginal of the state on a finite subset of the chain. We call this realization observable regular realization, as it is expressed only in terms of quantities which can be estimated from experiments.
@@ A.2.1 Observable regular realization
with $X\in\mathcal{C}_{R}$, has dimension $m$. We call this vector space $V$, which can be thought of as a subspace of ${\mathcal{A}_{\Lambda}^{*}\cong\mathcal{A}_{\Lambda}}$, for some finite subset $\Lambda$ of the chain, say $\Lambda=[1,t^{*}]$ for some $t^{*}\leq m$ (see Section A.2.2), and define $L$ to be a projection from $\mathcal{A}_{\Lambda}^{*}$ to $V$.
@@
Proposition A.5 (Observable regular realization).
@@
A.2.2 Guarantees on marginal size for translation invariant states
We observe that if the rank $m$ of the finitely correlated state viewed as a bilinear form is known, the construction in the previous subsection can be made completely explicit. First, we have the following lemma.
Lemma A.6.
Define $V_{i}$ the vector space generated by functionals $\Phi_{X}:\mathcal{A}_{R}\rightarrow\mathbb{C}$ of the form
$\Phi_{X}(A_{1}\otimes\cdots\otimes A_{n}):=\omega(X\otimes A_{1}\otimes\cdots\otimes A_{n}),$ (54)
with $X\in\mathcal{A}_{[-i+1,0]}$ and $n\in\mathbb{N}$. If $\omega$ has a minimal realization of dimension $m$, then there exists $i\leq m$ such that $V_{j}=V_{i}=\Phi(\mathcal{A}_{L})$ for each $j\geq i$. In particular, this holds for $V_{m}$.
Proof.
Let $i$ be such that $\dim V_{i}=\dim V_{i+1}$, i.e. $V_{i}=V_{i+1}$. Let $\Phi_{Y_{k}}$, $k=1,\dots,\dim V_{i}$, be a basis of $V_{i}$, so $Y_{k}\in\mathcal{A}_{[-i+1,0]}$. If $X=\sum_{j}X_{j}\otimes A_{j}\in\mathcal{A}_{[-i-1,-1]}\otimes\mathcal{A}=\mathcal{A}_{[-i-1,0]}$, i.e. $\Phi_{X}\in V_{i+2}$, then for $A\in\mathcal{A}_{R}$ using translation invariance of the FCS $\omega$:
$\Phi_{X}(A)=\sum_{j}\Phi_{X_{j}}(A_{j}\otimes A)=\sum_{j,k}c_{jk}\Phi_{Y_{k}}(A_{j}\otimes A)=\sum_{j,k}c_{jk}\Phi_{Y_{k}\otimes A_{j}}(A)=\Phi_{Y}(A)$
where $Y=\sum_{j,k}c_{jk}Y_{k}\otimes A_{j}\in\mathcal{A}_{[-i,0]}$. Therefore, $V_{i+2}=V_{i+1}$ and by induction $V_{i}=V_{j}=\Phi(\mathcal{A}_{L})$ for all $j\geq i$. Because of $\dim(\Phi(\mathcal{A}_{L}))=m$ by assumption, we have $\dim V_{i}=m$ and thus $i\leq m$ since $\dim V_{j}$ has to increase monotonically in $j$. ∎
Let $\Psi:\mathcal{A}_{R}\to\mathcal{A}_{L}^{*}$ be the map such that $\Psi_{X}(Y)=\Phi_{Y}(X)$ where $X\in\mathcal{A}_{R},Y\in\mathcal{A}_{L}$. We have $\Psi(\mathcal{A}_{R})\cong\Phi(\mathcal{A}_{L})^{*}$ since both are finite-dimensional of the same dimension and since $\Psi_{X}(\Phi_{Y})=\Phi_{Y}(X)$ is well-defined. By the above Lemma we can replace $\mathcal{A}_{L}$ by a finite subchain $\Psi(\mathcal{A}_{R})\cong\Phi(\mathcal{A}_{[-m+1,0]})^{*}$. The argument in the proof of the Lemma can also be applied to the analogous spaces $W_{i}$ for $\Psi$ showing that $\Psi(\mathcal{A}_{R})\cong\Psi(\mathcal{A}_{[1,m]})$. Therefore, it suffices to parametrize $\Psi_{X}$ by $X\in\mathcal{A}_{[1,m]}$ and restrict inputs to $\mathcal{A}_{[-m+1,0]}$ so one has:
Corollary A.7.
$\Psi_{X}(A):=\omega(X\otimes A),$ (55)
A.2.3 Quotient realizations
Let $(V,u,\mathbb{E},\sigma)$ be any realization and $(V^{\prime},e^{\prime},\mathbb{K}^{\prime},\rho^{\prime})$ its quotient realization. Since $(V^{\prime},e^{\prime},\mathbb{K}^{\prime},\rho^{\prime})$ and the observable realization $(\mathbb{C}^{m},e,\mathbb{K},\rho)$ constructed in Proposition A.5 are both regular, there is a unique invertible linear transformation $M:V^{\prime}\rightarrow\mathbb{C}^{m}$ such that
@@ A.2.3 Quotient realizations
A specific map $M$ with desirable properties can be obtained explicitly as follows: Recall that $\mathcal{C}_{R}\subseteq\mathcal{A}_{[1,t^{*}]}$ and $\mathcal{C}_{L}\subseteq\mathcal{A}_{[1-t^{*},0]}$ (there is always such $t^{*}$, in fact we can take $t^{*}\leq m$, see Appendix A.2.2). We define the map $\mathcal{F}:V\rightarrow\mathcal{A}_{[1-t^{*},0]}^{*}$ as
@@
Note, that $\mathcal{F}\mathbb{E}_{A}(u)=\Psi_{A}$ and thus $\ker(U^{\intercal})=\mathcal{F}(W)^{\perp}$. Now, $Z\in\ker(\mathcal{F})$ if and only if for each $A$ we have
@@
In particular, $(\mathbb{C}^{m},e,\mathbb{K},\rho)$ is similar to the quotient realization of $(V^{\prime},e^{\prime},\mathbb{E},\rho^{\prime})$ through the invertible map $M:V^{\prime}\rightarrow\mathbb{C}^{m}$ defined by $ML=U^{\intercal}\mathcal{F}|_{W}$ (see Proposition A.4).
@@
It holds that $\mathcal{F}(u)=\Omega(\mathds{1})$ so $U^{\intercal}\mathcal{F}(u)=e$ is immediate. Next, note that by definition $\rho U^{\intercal}\mathcal{F}=\tau\mathcal{F}$. Hence, if $w=\mathbb{E}_{A}(u)\in W$, $A\in\ker(\Omega)^{\perp}$, then
@@
Writing $U^{\intercal}\mathcal{F}=ML$ by Lemma A.8 the proof is complete. ∎
A.2.4 $C^{*}$-realizations
@@
We use the symbol $\mathcal{E}$ instead of $\mathbb{E}$ when dealing with $C^{*}$ realizations to distinguish them among realizations. Note, that whenever we find a realization with memory system being a $C^{*}$-algebra and generating map $\mathbb{E}$ being completely positive in the natural $C^{*}$ order, we can always find a realization satisfying the same that is of the form described in Definition A.10, even with $\rho_{0}$ being of full support, see [14, Lem. 2.5].
As before we will also use the notation $\mathcal{E}_{A}(B)=\mathcal{E}(A\otimes B)$. We can apply the construction of the quotient realization Proposition A.4, that is we set
@@
Lemma A.8 then takes the following form:
Lemma A.11.
@@
A.2.5 Empirical realizations
It will be useful to also consider the following realizations, obtained by a similarity transformation of the observable realization. In the following $\hat{U}$ can be thought as a truncated left orthogonal map in the SVD of $\hat{\Omega}$, which is an estimate of $\Omega$ (cf.  Appendix C). While the realization constructed from $\hat{\Omega}$ will not be an exact realization of $\omega$, the empirical realization constructed here is, and will be easier to compare with the estimate than the observable realization.
Proposition A.12 (Empirical realizations).
@@ A.2.5 Empirical realizations
Proof.
The proof is by explicit substitution, noting that $({\hat{U}}^{\intercal}\Omega)^{+}=({{U}}^{\intercal}\Omega)^{+}(\hat{U}^{\intercal}U)^{-1}$ and that $UU^{\intercal}\Omega_{A}=\Omega_{A}$ because the images of $\Omega$ and $\Omega_{(\cdot)}$ coincide. ∎
@@
where $M$ is the linear map connecting a quotient realization with the observable realization, as in Proposition A.9. The following proposition easily follows.
Proposition A.13 (Empirical realizations and quotient realization).
For an empirical realization as in Proposition A.12, and a quotient realization as in Proposition A.9, we have
@@
$\tilde{\mathbb{K}}_{A}\hat{M}L$ $=\hat{M}\mathbb{K}^{\prime}_{A}L=\hat{M}L\mathbb{E}_{A}|_{W},$ (78)
@@
Appendix B Operator systems and realizations
B.1 Operator systems and completely bounded norms
The memory system $V$ of a finitely correlated state $\omega$ on $\mathcal{A}$ with regular realization $(V,e,\mathbb{E},\rho)$ inherits the structure of an operator system [14, Lem. A.1]. In brief, it is constructed from the positivity structure on $\mathbb{M}_{n}(\mathcal{A})$ by push-forward with the map $\Psi:\mathcal{A}_{R}\to V$ and its amplifications.
Definition B.1.
  1. 1.
    Each $X\in C_{n}$ is self-adjoint.
  2. 2.
    $e$ is an Archimedean matrix order unit for the cones $C_{n}$.
  3. 3.
$S(X)=\begin{pmatrix}0&X\\ X^{*}&0\end{pmatrix}.$
$\|X\|_{e^{(n)}}=\inf\{r\in\mathbb{R}_{+}\ |\ re^{(2n)}\geq S(X)\}\geq 0.$
$e$ being Archimedean implies that the order norms are, in fact, norms and that the corresponding cones are closed. For self-adjoint $X$ the order norm is also given as $\|X\|_{e^{(n)}}=\inf\{r>0\ |\ re^{(n)}\geq X,-X\}$. It always holds that $\|e^{(n)}\|_{e^{(n)}}=1$. For a linear map $\alpha:V\to W$ between operator systems we denote its $n$-th amplification by $\alpha_{n}=\mathrm{id}_{\mathbb{M}_{n}}\otimes\ \alpha:\mathbb{M}_{n}(V)\to\mathbb{M}_{n}(W)$. $\alpha$ is called positive if it maps positive elements of $V$ to positive elements of $W$ and completely positive if each of its amplifications is positive. We then write $\alpha\in\mathrm{CP}(V,W)$. It is unital if it maps the order unit of $V$ to the order unit of $W$. If $\alpha$ is completely positive then $\alpha$ is bounded with its operator norm satisfying $\|\alpha\|=\|\alpha(e_{V})\|_{e_{W}}$. The Choi-Effros theorem [24, Thm. 13.1] ensures that every abstract operator system can be represented faithfully as a concrete operator system, meaning a *-invariant subspace of $\mathcal{B}(\mathcal{H})$ for some Hilbert space $\mathcal{H}$.
The estimates in Appendix C also require us to use norms of complete boundedness. The natural setting for talking about these norms is that of an operator space, that is a closed subspace of $\mathcal{B}(\mathcal{H})$. For more information see [44]. There is an abstract characterization of this structure, similar to the abstract characterization of operator systems above, such that every abstract operator space can be represented as a concrete operator space by Ruan’s theorem [24, Thm. 13.4]. The structure of an (abstract) operator system induces the structure of an (abstract) operator space. A linear map $\alpha:V\to W$ between operator spaces is called completely bounded if all its amplifications $\alpha_{n}$ are bounded and the sequence of operator norms $\|\alpha_{n}\|$ is uniformly bounded. The corresponding norm of complete boundedness is then
If $\alpha$ is a completely positive map between operator systems then $\|\alpha\|_{cb}=\|\alpha\|=\|\alpha(e_{V})\|_{e_{W}}$. If $\alpha$ is unital it is thus automatically completely contractive. Furthermore, the following bound is true [44, Thm. 3.8]:
Lemma B.2.
We also have the following result. It is basically [44, Prop. 1.12] which is valid also for abstract operator spaces by Ruan’s theorem.
Lemma B.3.
Remark B.4.
While there is a clear link between the structures of operator systems and operator spaces, the corresponding categories do not behave exactly the same. For example, the natural operator system structure induced on the quotient of an operator system does not in general coincide with the quotient operator space structure. A second subtlety concerns conventions in the literature. Concrete operator systems and, in particular, spaces are often assumed to be closed (hence complete) in norm. For the abstract structures, in particular operator systems, this is often not required (and then also not true when represented faithfully). This is certainly not a real problem in practice but should be kept in mind to avoid confusion. Both subtleties are illustrated by [45, Prop. 4.5].
Let $\mathcal{H}$ be a Hilbert space. Then there exists a natural operator space structure on $\mathcal{H}$ satisfying that the (conjugate linear) identification $\mathbb{M}_{n}(\mathcal{H})\cong\mathbb{M}_{n}(\mathcal{H})^{*}$ is completely isometric [44, Ch. 7]. We will implicitly use this operator system structure whenever we talk about complete boundedness of maps on Hilbert spaces. It holds [44, Prop. 7.2]:
Lemma B.5.
B.2 Operator systems and finitely correlated states